Let s= {p/q €Q|(q,7) =1} Define a relation ~ on s by p/q ~ a/b iff 7| (bp-aq). check whether or not ~ is an equivalence relation on s?
1
Expert's answer
2020-12-29T16:02:20-0500

The relation is an equivalence relation if:

  1. reflexitivity
  2. symmetry
  3. transitivity

Let

"s=\{\frac{p}{q}\in Q|(q,7)=1\}\\
\frac{p}{q}\sim\frac{a}{b} iff \frac{7}{bp-aq}"

Reflexitivity

For any fraction

"\frac{a}{b}, \frac{a}{b} \sim \frac{a}{b}\\
\frac{7}{ab-ab}=\frac{7}{0}" is not reflexitivity


 The relation ~ is not an equivalence relation on s